OpenStudy (anonymous):

Insert <, >, or = to make the sentence true. √3 √7 I'm thinking the answer is A A. = B. < C. >

jagr2713 (jagr2713):

3 is less than 7

jimthompson5910 (jim_thompson5910):

If \[\Large \sqrt{3} = \sqrt{7}\] then you could square both sides to get \[\Large \sqrt{3} = \sqrt{7}\] \[\Large (\sqrt{3})^2 = (\sqrt{7})^2\] \[\Large 3 = 7\]

jimthompson5910 (jim_thompson5910):

However, \[\Large 3 = 7\] is false so \[\Large \sqrt{3} = \sqrt{7}\] is false

OpenStudy (anonymous):

Ok thanks you guys so much!